DX AO tests on night of 20170205 UT

See Commissioning wiki for additional details Com20170205

03:59 UT

Loop closed everything seems stable.

Taking some Optical Loop Data.

20170205_040328 Data set no IRTC

20170205_040627 No IRTC (Not Working)

Optical Loop Data Corresponding to Vibration Tests/ IRTC data.

041805 10,000 frames

041959 10,000 frames

043122 10,000 frames

Doug's Notes

02:50 adsc_stop/start, stop/start wfsarb, wfs msgdrtdb died. Had to w_stop/start

03:05 WFS Operate: Power was already on, Main power on returns immediately, then get
Error
LOST COMMUNICATION WITH SLOPE COMPUTER

03:59 Opt Loop data
  • 035951

04:09 IRTC cube
  • exp = 0.01
  • frames 100
  • 040909.fits
  • 2.fits

04:18 cube
  • exp = 0.01
  • frames = 1000
  • 041806.fits
  • 3.fits

04:20 cube
  • exp = 0.01
  • frames = 10000
  • 042002.fits
  • 4.fits

04:30 Opt Loop Data
  • 043014

04:35 cube
  • open loop data
  • exp = 0.01
  • frames = 1000
  • 043513
  • 5.fits

04:39 PresetAO failed on DX. msg deamon die and many other processes

04:40 cube
  • dark
  • exp 0.01
  • frames = 1000
  • 044005.fits
  • 6.fits

04:44 PresetAO fail. need to check pointing

04:54 left RunAO alert for LUCI
  • RunAO command failed
  • WFS arbitrator find delta
    • x=363.73, Y=532.12
    • error X=-1.401 mm, Y=0249 mm
    • arbFunc doCheckRef done
    • No error message of failure
  • AOS gave AcquireRef{FLAO} Mode: Repoint telescope , Automatic acquire
    • No error message of failure

-- %USERSIG{GregoryTaylor - 2017-02-05}%
Topic revision: r4 - 20 Feb 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback